# Simulating Dual-Pixel Images From Ray Tracing For Depth Estimation

### v1

@@ 8.1 Structure parameters of DP composite pixels
To demonstrate the similarity between real and simulated DP PSFs under the optimal combination of structure parameters, we conduct a qualitative analysis while keeping the lens, F-number, kernel size, and focal distance unchanged from the search experiments. As shown in Fig. 12, we select five points at equal intervals within the valid imaging region, with only their x-coordinates increasing sequentially. Additionally, to visually compare changes in pixel values, a line plot of pixel distribution along the rows is also included in Fig. 12. Observing the actual DP PSF, we notice that as the object point $p$ moves farther from the optical axis, the real $PSF_{L}$ and $PSF_{R}$ become more asymmetrical. This observation aligns with the experimental results presented in Sec. 5.3 (Evaluation on DP PSF simulator) of the main paper. When comparing the ray-traced and real DP PSF, we find that the ray-traced DP PSF is not only highly realistic in morphology but also closely matches the real data in pixel values.
8.2 Local DP PSF convolution
@@
Use the different DP PSF kernel for different pixels.
@@
8.3 DP PSF predict network
Figure 12: DP PSF predict network representation result. Evaluate the ray-traced and network-predicted (ours, Yang et al. [39], and Tseng et al. [33]) F/4 DP PSFs at three depths (0.5m, 1.5m, 20m) and positions. For small DP PSF radii, all network predictions closely match ray-traced results. For larger radii, the predicted results by [39, 33] show significant deviations, attributed to differences in normalization schemes with ours during training.
@@ 8.3 DP PSF predict network
We present qualitative and quantitative comparison results between our MLP network and recent work [39, 33], As shown in Fig. 12, we evaluate DP PSF at three depths and positions, using ray-traced DP PSF as GT. Their results show significant errors compared to ray-traced ones at large PSF radii. This is due to their training scheme using sum normalization for GT, which is effective for small PSF radii but becomes less accurate for larger ones. In contrast, our network, also using only MLP, can predict accurate results at all depths and positions, even at large PSF radii. The only difference is that we adopt max normalization for GT DP PSF during training and mimic vignetting correction to scale predicted DP PSF when inference. Compared to the DP PSF obtained through ray tracing, the results output by the MLP network are smoother. This is because the number of rays sampled in ray tracing is limited by memory and computation time, with only 4096 rays set.
Table 6: Quantitative results of DP PSF predict networks.
@@
We present quantitative results in Tab. 6. Specifically, we compare L1 and L2 errors for 50 different depths and positions, as well as the time cost for predicting a DP PSF map for all pixels (a total of $512\times 768$) in the depth map. Our network outperforms those developed by Yang et al. [39] and Tseng et al. [33] in accuracy, and its low time cost enables high-speed image rendering. In contrast, the network proposed by Tseng et al. [33] incurs high time costs due to its inclusion of convolutional layers.
@@
9 Depth-from-Dual-Pixel model details
9.1 Adjustment of cost volume
In the field of Depth-from-Dual-Pixel (DfDP), we do not pursue the optimal DfDP network structure. Instead, we select [6] as the DfDP model and make reasonable adjustments to its cost volume step to accommodate the DP data.
Existing binocular disparity matching algorithms typically only consider a single-direction misalignment to stack left and right image features, forming a cost volume. This is because in binocular images, the disparity formed by points at any depth is always in the same direction. However, in DP images, the disparity direction formed by object points before and after the focal distance is opposite. Therefore, as shown in Fig. 4(c) (main paper), we add reverse disparity while stacking the original disparity. Our implementation is as follows:
@@ 9.2 A new test set DP119
We used the Canon RF50mm lens and Canon R6 Mark II camera to capture the dataset, setting the aperture to F/4 and focusing at 1 meter. Each scene provides real DP-depth paired data. Finally, we collected the DP119 dataset, which includes 45 planar scenes, 44 box scenes, and 30 casual scenes, totaling 119 scenes.
@@
Figure 13: Examples of scenes in the DP119 dataset. Each depth map is decorated with a color bar in meters. (a) Planar scene samples. The d is the planar depth value at the time of capture. (b) Box scene samples. (c) Casual scene samples.
@@
For the planar scenes, we captured richly textured posters images from 0.5 meters to 2 meters perpendicular to the camera’s optical axis, with denser sampling at shallower depths of field. The dataset includes five sets of samples and nine different distances (0.5m, 0.6m, 0.7m, 0.8m, 0.9m, 1.0m, 1.2m, 1.5m, 2m). Real DP images were captured at apertures F/4 and F/20, adjusting ISO to match light intensity. F/20 images were used as AiF images, with their depth value copied into the depth map. Due to the simple depth structure of planar scenes, we avoid issues like holes in LiDAR depth maps or misalignment between depth and RGB images. These scenes are useful for evaluating both depth estimation models and the realism of DP simulators. As shown in Fig. 13(a), we present a sample set of planar scenes.
@@
For the box scenes, we covered all the boxes and backgrounds with posters. The boxes were randomly placed within a range of 0.5m to 2m, and each time we took a shot, we significantly adjusted the number, texture, and placement of the boxes, deliberately creating situations of overlap, tilt, and occlusion. Depth was captured using the LiDAR of an iPhone 15 Pro. Both planar and box scenes are richly textured, aiding the model in utilizing aberration and phase difference cues for depth estimation. Textureless areas, even when defocused, do not introduce any aberrations or phase differences, which would interfere with depth estimation. Therefore, planar and box scenes are ideal for evaluating whether the DP simulator addresses the domain gap between simulated and real DP images. As shown in Fig. 13(b), we present samples of box scenes.
@@
For the casual scenes, we directly captured ordinary real-world scenes, intentionally controlling the depth within 0.5m to 10m, as scenes beyond 10m would yield unreliable results from the depth sensor. Among the 30 casual scenes, 20 indoor scenes were captured with depth obtained using a binocular structured-light camera (Orbbec Gemini2). For the 10 outdoor scenes, since the binocular structured-light camera performs poorly outdoors, we used the LiDAR of an iPhone 15 Pro to capture depth. As shown in Fig. 13(c), we present samples of casual scenes.
9.3 More DfDP results on the DP119 test set
@@
Figure 14: The depth estimation results from different simulators on the DP119 test set. Evaluate DfDP models by with L2R [2], Modeling [26], DDDNet [23] and our Sdirt on various DP119 dataset scenes. Each result image is decorated with a color bar in meters. Their depth estimation results show partial accuracy in relative positional relationships but large absolute positional errors. Our depth estimation results, however, demonstrate accurate in both relative and absolute positions with minimal errors. Furthermore, textureless areas lead to degradation in all models. (Best viewed in colour and enlarge on screen.)
@@
Fig. 14 presents the depth estimation results of various simulators on different scenes in the DP119 dataset. From planar scenes 1 and 2, we observe that due to increased aberrations and asymmetric phase differences at the edge regions, all models, including ours, exhibit significant relative positional errors in both central and edge regions. However, our model provides more accurate absolute depth estimates, whereas other models tend to bias toward the focus distance (1m). The depth estimation results of L2R [2] and Modeling [26] are similar, as their DP PSF peak positions are spatially close. A similar trend is observed in box scenes 1 and 2, where other models show a strong bias toward the focal distance (1m) and exhibit large relative positional errors in both central and edge regions. In contrast, our model maintains superior performance.
@@
In casual scenes 1 and 2, the textures are rich, and the scenes are simple. In addition to aberration and phase difference information, the models can also use additional structural information to predict depth. As a result, the relative positional errors in the center and edge regions were alleviated for all models in these two scenes. However, the absolute depth estimates from other models still show a clear bias toward the focus distance (1m). Moreover, in casual scenes 3 and 4, large textureless areas appear, missing aberration and phase difference cues. Depth estimation models can only infer the depth of these textureless areas along structural information, leading to significant relative positional position errors in textureless regions for all models.

### v2

@@ Acknowledgments
To demonstrate the similarity between real and simulated DP PSFs under the optimal combination of structural parameters, we conduct a qualitative analysis while keeping the lens, F-number, kernel size, and focal distance unchanged from the grid search experiments. As shown in Fig.˜S3, we select five points evenly spaced within the valid imaging region, with only their x-coordinates increasing sequentially. Additionally, to visually compare changes in pixel values, a line plot of pixel distribution along the central row is also included in each subplot. Observing the actual DP PSF, we notice that as the object point $p$ moves farther from the optical axis, the real $PSF_{L}$ and $PSF_{R}$ become more phase asymmetric. This observation aligns with the experimental results presented in Sec. 5.3 (Evaluation on simulated DP PSFs) of the main paper. When comparing the ray-traced and real DP PSFs, we find that the ray-traced DP PSF is not only highly realistic in morphology but also closely matches the real data in pixel values.
S1.2 DP PSF capture and linearization
The native resolution of the Canon R6 Mark II is $6000\times 4000$ pixels. Due to GPU memory limitations, we reduced the spatial resolution to $768\times 512$. Although this resolution is far below the original, our experiments on DP PSFs (see Fig. 5 in the main paper and Fig.˜S3) demonstrate that aberrations and phase differences in defocused regions remain clearly visible.
Capturing the real DP PSFs is a straightforward process. A single white pixel is displayed on an OLED screen (iPhone 15 Pro, 460 ppi, 55 um pixel size). We capture the PSFs using a Canon R6 Mark II (6 um pixel size) equipped with an RF50mm lens at object distances ranging from 0.5 m to 2 m. For the camera’s native resolution of $6000\times 4000$, the Nyquist-limited spatial resolution is approximately 120 um at 0.5 m and 480 um at 2 m. Given the subsequent downsampling, capturing a single bright pixel on the screen is sufficient to ensure the acquisition of the real PSFs.
The PSFs generated by optical simulations represent linear light intensity distributions, similar to those in the camera’s RAW domain. To our knowledge, Canon’s RAW-domain dual-pixel images are not publicly available, and currently only RGB-domain dual-pixel images can be obtained. Therefore, the PSFs captured by the camera are pixel value distributions in the RGB domain, transformed from RAW-domain light intensities by the image signal processing (ISP) pipeline. This pipeline involves non-linear and non-invertible operations such as gamma correction and contrast enhancement. Although modern ISP pipelines are highly complex, radiometric calibration remains a common and practical approach to approximately linearize RGB data and recover the underlying light intensity distributions in the RAW domain. As shown in Fig.˜S4, our linear calibration method consists of: 1. Capture a static scene under varying exposure times; 2. Use the image captured with the shortest exposure time (pixel values within [0, 35]) as a linear reference, scaling it according to exposure time ratios to obtain target values for the other images; 3. Fit the original pixel values of other images to these targets. We adopt a piecewise weighted fitting approach, and the fitted functions are released with our code.
Figure S4: Overview of our linear calibration method. The plot shows three components: pixel values from images captured under varying exposure times, target values derived by scaling the shortest exposure image (pixel values within [0, 35]) according to exposure time ratios, and the fitted curve obtained by applying a piecewise weighted fitting approach to align original pixel values with target values. The fitted functions are publicly available in our released code.
S1.3 Local DP PSF convolution
@@
Use different DP PSFs for different pixels.
@@
S1.4 DP PSF prediction network
Figure S5: Qualitative results of DP PSF prediction networks. Evaluate the ray-traced and network-predicted (ours, Yang et al. [39], and Tseng et al. [33]) F/4 DP PSFs at three depths (0.5m, 1.5m, 20m) and positions. For small DP PSF radii, all network predictions closely match ray-traced results. For larger radii, the predicted results by [39, 33] show significant deviations, attributed to differences in normalization schemes with ours during training.
@@
We present qualitative and quantitative comparison results between our MLP network and recent work [39, 33], As shown in Fig.˜S5, we evaluate DP PSF at three depths and positions, using ray-traced DP PSF as GT. Their results show significant errors compared to ray-traced ones at large PSF radii. This is due to their training scheme using sum normalization for GT, which is effective for small PSF radii but becomes less accurate for larger ones. In contrast, our network, also using only MLP, can predict accurate results at all depths and positions, even at large PSF radii. The only difference is that we use max normalization for GT DP PSFs during training, and apply sum normalization to the predicted DP PSFs during inference to approximate the uniform intensity distribution of cameras with vignetting compensation. Compared to the DP PSF obtained through ray tracing, the results output by the MLP network are smoother. This is because the number of rays sampled in ray tracing is limited by memory and computation time, with only 4096 rays set.
Table S3: Quantitative results of DP PSF prediction networks.
@@
We present quantitative results in Tab.˜S3. Specifically, we compare L1 and L2 errors for 50 different depths and positions, as well as the time cost for predicting a DP PSF map for all pixels (a total of $512\times 768$) in the depth map. Our network outperforms those developed by Yang et al. [39] and Tseng et al. [33] in accuracy, and its low time cost enables high-speed image rendering. In contrast, the network proposed by Tseng et al. [33] incurs high time costs due to its inclusion of convolutional layers.
@@
S2 Depth-from-Dual-Pixel Details
S2.1 Adjustment of cost volume
In the field of DfDP, we do not pursue the optimal DfDP network structure. Instead, we select [6] as the DfDP model and make reasonable adjustments to its cost volume step to accommodate the DP data.
Existing binocular disparity matching algorithms typically only consider unidirectional disparity shifts when stacking left and right image features to form the cost volume. This is because in binocular images, the disparity formed by points at any depth is always in the same direction. However, in DP images, the disparity direction formed by object points before and after the focal distance is opposite. Therefore, as shown in Fig. 4(d) (main paper), we add reverse disparity while stacking the original disparity. Our implementation is as follows:
@@
We used the Canon RF50mm lens and Canon R6 Mark II camera to capture a new dataset, setting the aperture to F/4 and focusing at 1 m. Each scene provides real DP-depth paired data. Finally, we collected the DP119 dataset, which includes 45 planar scenes, 44 box scenes, and 30 casual scenes, totaling 119 scenes.
@@
Figure S6: Examples of scenes in the DP119 dataset. Each depth map is decorated with a color bar in meters. (a) Planar scene samples. The $d$ is the planar depth value at the time of capture. (b) Box scene samples. (c) Casual scene samples.
@@
For the planar scenes, we captured richly textured posters images from 0.5 m to 2 m perpendicular to the camera’s optical axis, with denser sampling at shallower depths. The dataset includes five distinct scenes at nine different distances (0.5 m, 0.6 m, 0.7 m, 0.8 m, 0.9 m, 1.0 m, 1.2 m, 1.5 m, 2 m). Real DP images were captured at apertures F/4 and F/20, adjusting ISO to match light intensity. F/20 images were used as AiF images, with their depth value copied into the depth map. Unlike box and casual scenes, planar scenes have simple depth structures, which avoid common issues such as holes in LiDAR depth maps and misalignment between depth and RGB images. These scenes are useful for evaluating both depth estimation models and the realism of DP simulators. As shown in Fig.˜S6(a), we present a sample set of planar scenes.
@@
For the box scenes, we covered all the boxes and backgrounds with posters. The boxes were randomly placed within a range of 0.5 m to 2 m, and each time we took a shot, we significantly adjusted the number, texture, and placement of the boxes, deliberately creating situations of overlap, tilt, and occlusion. Depth was captured using the LiDAR of an iPhone 15 Pro. Both planar and box scenes are richly textured, aiding the model in utilizing aberration and phase difference cues for depth estimation. Textureless areas, even when defocused, do not introduce any aberrations or phase differences, which would interfere with depth estimation. Therefore, planar and box scenes are ideal for evaluating whether the DP simulator addresses the domain gap between simulated and real DP images. As shown in Fig.˜S6(b), we present samples of box scenes.
@@
For the casual scenes, we directly captured ordinary real-world scenes, intentionally controlling the depth within 0.5 m to 10 m, as scenes beyond 10m would yield unreliable results from the depth sensor. Among the 30 casual scenes, 20 indoor scenes were captured with depth obtained using a binocular structured-light camera (Orbbec Gemini2). For the 10 outdoor scenes, since the binocular structured-light camera performs poorly outdoors, we used the LiDAR of an iPhone 15 Pro to capture depth. As shown in Fig.˜S6(c), we present samples of casual scenes.
S2.3 More DfDP results on the DP119 test set
@@
Figure S7: The depth estimation results from different simulators on the DP119 test set. Evaluate DfDP models with L2R [3], Modeling [25], DDDNet [23] and our Sdirt on various DP119 dataset scenes. Each result image is decorated with a color bar in meters. Their depth estimation results show partial accuracy in relative positional relationships but large absolute positional errors. Our depth estimation results, however, demonstrate accurate in both relative and absolute positions with minimal errors. Furthermore, textureless areas lead to degradation in all models. (Best viewed in color and enlarge on screen.)
@@
Fig.˜S7 presents the depth estimation results of various simulators on different scenes in the DP119 dataset. From planar scenes 1 and 2, we observe that due to increased aberrations and asymmetric phase differences at the edge regions, all models, including ours, exhibit significant relative positional errors in both central and edge regions. However, our model provides more accurate absolute depth estimates, whereas other models tend to bias toward the focus distance (1 m). The depth estimation results of L2R [3] and Modeling [25] are similar, as their DP PSF peak positions are spatially close. A similar trend is observed in box scenes 1 and 2, where other models show a strong bias toward the focal distance (1 m) and exhibit large relative positional errors in both central and edge regions. In contrast, our model maintains superior performance.
@@
In casual scenes 1 and 2, the textures are rich, and the scenes are simple. Models can predict depth not only using aberrations and phase differences but also by leveraging additional structural cues. As a result, the relative positional errors in the center and edge regions were alleviated for all models in these two scenes. However, the absolute depth estimates from other models still show a clear bias toward the focus distance (1 m). Moreover, in casual scenes 3 and 4, large textureless areas appear, missing aberrations and phase differences. Depth estimation models can only infer the depth of these textureless areas along structural cues, leading to significant relative positional errors in textureless regions for all models.